OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (3, 6) and (–2, 1)?

OpenStudy (anonymous):

First use the slope formula to find the slope. \[m = \frac{y_1 - y_2}{x_1 - x_2} \implies \frac{6 - 1}{3 - (-2)} \implies ?\]Then use point-slope form. \[y - y_1 = m(x - x_1)\]

OpenStudy (zzr0ck3r):

or you can use two equations 6 = 3m+b 1 = 2m+b multiply second by -1 6=3m+b -1 = -2m-b add them 5 = m if m = 5 6 = 3(5) +b b = -9

OpenStudy (zzr0ck3r):

y = 5x-9

OpenStudy (anonymous):

@zzr0ck3r It's asking for point-slope form. Not slope-intercept form.

OpenStudy (zzr0ck3r):

o woops ty sir

OpenStudy (zzr0ck3r):

ignore me then, you can get thier but not much fun:)